OpenStudy (anonymous):

Find (f+g)(x) (f-g)(x) (fg)(x) (f/g)(x) what is the domain of f/g? 1.) f(x)=x^2+3x+2 g(x)=x

OpenStudy (angela210793):

f+g means u add both eq and since u were laughing with Neillene's q i guess u know how to add them :)

OpenStudy (anonymous):

f/g(x) = x + 3 + 2/x so the domain is where this function is defined It has a singularity at x=0 so the domain is all real numbers except x=0

OpenStudy (anonymous):

(f+g)(x) = x^2+4x+2 (f-g)(x) = x^2+2x+2 (fg)(x) = x^3+3x+2x

OpenStudy (angela210793):

(fg)(x) = x^3+3x^2+2x

OpenStudy (anonymous):

thank you so much guys! especially angela210793.. ;D